OpenStudy (anonymous):

Solve for d. 4d + 7 = -17 d = -6 d = -2.5 d = 2.5 d = -4.25

OpenStudy (nirmalnema):

4d + 7 = -17 4d=-17-7 4d=-24 d=-24/4 d=-6

OpenStudy (anonymous):

Whaa??

OpenStudy (nirmalnema):

???

OpenStudy (anonymous):

nevermind i got the answer.

OpenStudy (nirmalnema):

ok.